The demand for money function was MD = 4Y − 1000i where MD is the quantity of money demanded, i is the rate of interest (interest of 5 means 5 percent in this problem), and Y is real national income, which currently is 1500. The supply of money is 1000, currency in circulation outside the banking system is 100, the target reserve ratio is 10 percent, there is no cash drain in the banking system, and the recessionary gap is 250. The price level does not change


1.  Suppose the Bank of Canada reduces the target for the overnight rate, prompting commercial banks to reduce the market rate of interest to 4 percent. Are the commercial banks experiencing a situation of excess cash reserves or of too little cash reserves? What is the size of this excess/insufficient cash reserve when Y = 1500?



2.  What will the commercial banks do to eliminate excess/insufficient cash reserves? By how much should the level of cash reserves of the banking system change for the economy to move to full employment?




1
Expert's answer
2021-04-06T06:59:45-0400
"solution"

A]

"Md=4\times 1500-1000i"

"Md=6000-1000i"

"Md=6000-1000[{4}]"

"Md=6000-4000"

"Md=2000"

The commercial banks experiencing a situation of excess cash reserves

"2000-1500=500"


B]The central bank can also use the IOER rate as a tool of monetary policy. By raising the IOER rate, the central bank gives commercial banks more incentives to hold excess reserves, which reduces the money supply. To conduct an expansionary monetary policy, the central bank can lower the IOER rate. This will lead to commercial banks reducing their excess reserves.

it should reduce by 500